# Predictive Traffic Rule Compliance using Reinforcement Learning

### v1

@@ IV Experiments
This chapter details our experimental investigation, spanning the foundational setup, implementation details, and a comparative analysis of our modified reinforcement learning (RL) approach against a baseline model. We leverage the highD dataset for evaluation. All experiments were conducted on a computing platform equipped with 12 CPUs, an Nvidia RTX 2060 GPU, and 30 GB of RAM. The training process was organized into three distinct phases, each targeting a specific traffic rule and comprising 20,000 steps.
@@ IV-A Dataset
To adapt the dataset for our framework, we utilized the CommonRoad-Dataset Converter [18] to transform highD data into the CommonRoad format. During conversion, we adjusted the temporal resolution to 0.1 s intervals, optimizing the data structure for our learning algorithm while retaining sufficient granularity for precise trajectory analysis.
@@ IV-B Experimental Setup
The comprehensive hyperparameters for the environment, model, and training are presented in Table III.
@@
RL Actor-critic networks $(\pi,V)$ Each $MLP(256,128,64)$
@@
The experimental environment is designed with specific spatial configurations to ensure effective training scenarios. Each scenario includes a goal region randomly positioned 10 m to 50 m before the road’s terminus. The ego vehicle is initialized 150 m to 350 m from the goal, with an initial velocity of 15 m/s and a minimum clearance of 3 m from surrounding obstacles. This minimum distance allows the ego vehicle to spawn in challenging positions, enabling exploration of complex states.
At each training iteration, 256-step rollouts are collected and the return is computed using a discount factor of 0.99. This high value emphasizes the avoidance of long-term rule violations, prioritizing sustained reward accumulation over short-term gains.
@@
After each rollout step, the critic network is updated using collected states and rewards derived from the robustness of the rules with hyperparameters defined in Table III. The reactive planner’s cost function integrates these updated value estimates from the critic network, enabling continuous interaction with the environment through on-policy learning.
@@ IV-C Rewards Design
We implement an independent training phase for each traffic rule, employing rule-specific robustness measures as reward signals with the same features, agent, and model. This separation facilitates focused learning of individual rule characteristics, minimizing cross-interference and eliminating the need to define and tune weights prior to training, thereby significantly streamlining the process.
Despite robustness-based rewards, we identified a challenge due to the absence of a termination reward. Without a positive reward for reaching the destination, and with a sparse or zero reward function throughout an episode, the cumulative future reward (i.e., the return) diminishes as the ego vehicle approaches the goal. This reduction occurs because fewer remaining steps limit opportunities to accumulate rewards, resulting in a counterintuitive decrease in the state-value function $V(s)$ near the destination—contrary to the expectation that states closer to the goal should have higher value due to their desirability. To address this, we introduce a TrajectoryProgressionRewardComputer, which provides positive rewards for advancing toward the goal, ensuring that the agent learns a policy that efficiently prioritizes its achievement.
@@
TABLE IV: Reward computer for reinforcement learning
@@ IV-D Benchmarking
To assess our approach, we established a non-learning baseline motion planner, employing the same reactive planner parameters as our method but incorporating traffic rule robustness costs directly. This baseline evaluates the trajectories by explicitly integrating compliance measures. For comparison, we defined a baseline cost function that substitutes learned value predictions with immediate rule evaluations via online robustness monitoring, which fundamentally differs from our learning-based approach, as it is limited to evaluating the immediate planning horizon. Consequently, it lacks the predictive capacity to foresee and mitigate rule violations beyond the current window, a limitation that our method overcomes through RL-driven foresight.
@@ IV-E1 Explained Variance
The explained variance measures how well the critic value function $V_{\phi}(s)$ predicts the actual returns (cumulative rewards) observed during training. It is defined as:
@@
where $G_{t}$ is the true return (discounted sum of rewards from timestep $t$), $V_{\phi}(s_{t})$ is the predicted value, and Var denotes variance. A value close to 1 indicates that the critic accurately captures the variability in returns, while a value close to 0 or negative suggests poor prediction.
@@
In Figure 3, the explained variance graph illustrates distinct learning dynamics for the traffic rules over 400 training steps. Specifically, $R\_{G1}$ demonstrates a steady increase, rising from 0.0 to approximately 0.9. This consistent ascent is particularly pronounced up to step 200, after which it stabilizes with minor fluctuations, reflecting the model’s progressively improving ability to predict returns. In contrast, $R\_{J2}$ exhibits greater volatility, characterized by notable peaks and troughs, while $R\_{J6}$ shows a slower, more gradual increase. The steady progress of $R\_{G1}$ indicates that the model effectively enhances its capacity to ensure long-term compliance with the safe distance rule, aligning with our evaluation objectives. Additionally, $R\_{I6}$ emerges as the easiest rule to learn, as it involves no vehicle-to-vehicle interactions, whereas $R\_{I2}$ proves the most challenging due to the extensive vehicle interactions it requires the model to capture.
@@ IV-E2 Episode Reward Mean
The episode reward mean is the average total reward collected over an episode, a complete driving scenario from start to finish. For our setup, this likely reflects the cumulative robustness or compliance with traffic rules like $R_{G1}$, averaged across multiple episodes:
@@
The Episode Reward Mean Comparison graph evaluates three configurations over 400 steps, with rewards ranging from 0 to 300. The blue line ($R\_{GI_{1}}$) starts at $\sim$300 but drops to $\sim$150 by step 50, stabilizing thereafter, possibly due to early overfitting. The red line ($R\_{GI_{2}}$) begins at $\sim$50, peaks at $\sim$200, and fluctuates between 150-250, indicating inconsistent performance. The green line ($R\_{16}$) starts near 0 and rises steadily to $\sim$300 by step 400, surpassing the others with a robust upward trend. This suggests that $R\_{16}$ achieves the highest long-term reward, reflecting better adaptability and learning efficiency.
@@ IV-E3 Performance of R_I6 Predictive Model
Our primary focus is on the custom rule $R_{I6}$, for which we compare the trained value network with the baseline model within the same scenario, featuring a no overtaking sign positioned at 300m along the longitudinal axis. We generate heatmaps using a grid-segmented representation of the road, evaluating the approximated state-value and $R\_{I6}$ rule robustness for both models. As depicted in Figure 5, a yellow grid indicates higher values, while a purple grid signifies lower values. The figure presents a comparative analysis of the trained value network and the baseline model, utilizing heatmaps overlaid on the road grid to assess the approximated state-value (top heatmap) and $R_{I6}$ rule robustness (bottom heatmap). The color gradient spans from purple (low values) to yellow (high values), with blue vehicles marking the ego vehicle’s positions and gray dashed lines indicating lane markings or potential trajectories.
Baseline Model: The bottom heatmap displays a sharp yellow-to-purple transition between 250m and 300m, attributable to the traffic sign detection range being set at 50 m. This transition forms a rectangular shape, as the model increasingly avoids leftward positions as it approaches the sign. This suggests that the baseline model relies heavily on immediate rule evaluation, with limited foresight into the no overtaking constraint.
Trained Value Network: The top heatmap exhibits a smooth transition from yellow to purple, approximately aligning with the baseline model. However, the initial value change begins earlier, around 230m, compared to the 250m detection range, the range increased from 50m to around 70m, indicating an awareness of long-term rule violation risks. Notably, the heatmap adopts a trapezoidal shape, with the value in the left lane decreasing earlier than in the right lane. Leveraging the predictive model, the lateral position of the ego vehicle influences the timing of lane-changing maneuvers: a vehicle positioned farther to the left can initiate a rightward lane change earlier, whereas a vehicle already near the right lane may delay the maneuver without violating the rule. This behavior aligns with human intuition and enhances safety and efficiency, as it enables proactive adjustments tailored to the vehicle’s spatial context.
@@ IV-F Discussion
This experiment demonstrates the model’s capability to predict long-term violations of traffic rules within the highD dataset scenarios. The proposed RL approach facilitates safer and more efficient trajectory planning by enabling the system to anticipate potential rule violations and adjust maneuvers proactively, reducing the likelihood of entering states that lead to unavoidable violations. For rules $R_{I6}$ and $R_{G1}$, which involve constraints independent of direct vehicle-to-vehicle interactions or focus on maintaining safe distances, the method exhibits strong performance, as evidenced by the smoother transitions in the value network outputs and the expanded safety regions in the heatmaps.
However, the model shows reduced robustness for rules involving complex vehicle interactions, such as $R_{I2}$, which requires maintaining appropriate speeds relative to surrounding vehicles. This limitation is reflected in the volatility observed in the explained variance metric and the inconsistent value predictions in the heatmap analysis. The challenge arises from the intricate dynamics of vehicle interactions, which demand a more detailed understanding of relative positions, speeds, and potential maneuvers. Addressing this issue may require enhancements such as increasing the volume of training data, extending the training duration, and incorporating additional features, such as relative velocity or lane-changing probabilities, to better capture the interaction dynamics. Furthermore, adopting a more advanced scenario extraction model, such as a Heterogeneous Graph Neural Network, could enhance the model’s ability to represent and manage complex traffic interactions by accounting for diverse vehicle types and their interconnected behaviors.
@@ V Conclusion
This thesis introduces a hybrid framework aimed at enhancing traffic rule compliance in autonomous driving systems. The approach addresses the limitation of short prediction horizons in traditional planning methods by integrating an actor-critic deep reinforcement learning (DRL) algorithm with a cost-based motion planner. This combination enables improved predictive compliance with traffic rules, particularly in complex highway scenarios on German interstate roads, as evaluated using the highD dataset. A Graph Neural Network (GNN) is utilized to construct a detailed state representation, allowing the critic network to learn state-value functions that anticipate potential rule violations beyond the immediate planning horizon. Experiments conducted on the highD dataset demonstrate that this method achieves better performance compared to a non-learning baseline, as evidenced by smoother transitions in state-value and robustness heatmaps for specific traffic rules, such as $R\_{I6}$ and $R\_{G1}$.
The proposed framework contributes to autonomous driving technology by developing a hierarchically structured rule book based on German traffic regulations and implementing a modular system compatible with cost-based motion planners. This modular design ensures flexibility, allowing the system to be adapted to various planning architectures with minimal adjustments.
In summary, this thesis provides a robust and adaptable solution for predictive traffic rule compliance in autonomous driving. By combining learning-based prediction with planning, the hybrid approach enhances compliance with traffic rules and supports safer navigation in dynamic environments. The framework establishes a foundation for future research in autonomous vehicle technology, particularly in improving the handling of complex vehicle interactions. As autonomous driving continues to develop, such frameworks may contribute to the realization of fully autonomous, rule-compliant systems, potentially leading to safer and more efficient roadways.

### v2

@@ IV Experiments
This chapter details our experimental investigation, spanning the foundational setup, implementation details, and a comparative analysis of our modified RL approach against a baseline model. We leverage the highD dataset for evaluation. All experiments were conducted on a computing platform equipped with 12 CPUs, an Nvidia RTX 2060 GPU, and 30 GB of RAM. The training process was organized into three distinct phases, each targeting a specific traffic rule and comprising 20,000 steps.
@@ IV-B Experimental Setup
The hyperparameters for the environment, model, and training are presented in Table III.
@@
AC networks $(\pi,V)$ Each $MLP(256,128,64)$
@@ IV-C Rewards Design
We implement an independent training phase for each traffic rule, employing rule robustness measures as reward signals with the same features, agent, and model. This separation facilitates focused learning of individual rule characteristics, minimizing cross-interference and eliminating the need to define and tune weights before training, thereby streamlining the process.
@@
TABLE IV: Reward computer for R
@@ IV-D Benchmarking
To assess our approach, we established a non-learning baseline motion planner. This baseline substitutes learned value predictions with immediate rule evaluations via online robustness monitoring. Consequently, it lacks the predictive capacity to foresee and mitigate rule violations beyond the current window.
@@ IV-E1 Explained Variance
The explained variance measures how well the critic value function $V_{\phi}(s)$ predicts the actual returns observed during training. It is defined as:
@@
where $G_{t}$ is the true return, $V_{\phi}(s_{t})$ is the predicted value, and Var denotes variance. A value close to 1 indicates that the critic accurately captures the variability in returns, while a value close to 0 or negative suggests poor prediction.
@@
In Figure 3, the explained variance graph illustrates distinct learning dynamics for the traffic rules over 400 training steps. Specifically, $R\_{G1}$ demonstrates a steady increase, rising from 0 to approximately 0.9. This consistent ascent is particularly pronounced up to step 200, after which it stabilizes with minor fluctuations, reflecting the model’s progressively improving ability to predict returns. In contrast, $R\_{J2}$ exhibits greater volatility, characterized by notable peaks and troughs, while $R\_{J6}$ shows a slower, more gradual increase. The steady progress of $R\_{G1}$ indicates that the model effectively enhances its capacity to ensure long-term compliance with the safe distance rule, aligning with our evaluation objectives. Additionally, $R\_{I6}$ emerges as the easiest rule to learn, as it involves no vehicle-to-vehicle interactions, whereas $R\_{I2}$ proves the most challenging due to the extensive vehicle interactions it requires the model to capture.
@@ IV-E2 Episode Reward Mean
The episode reward mean is the average total reward collected over an episode, a complete driving scenario from start to finish. For our setup, this likely reflects the cumulative robustness or compliance with traffic rules like $R\_{G1}$, averaged across multiple episodes:
@@
The Episode Reward Mean Comparison graph evaluates three configurations over 400 steps, with rewards ranging from 0 to 300. The $R\_{G1}$ starts at around 300 but drops to 150 by step 50, stabilizing thereafter, possibly due to early overfitting. The $R\_{G2}$ begins at 50, peaks at 250 by step 300, and fluctuates between 150-250, indicating acceptable training process. The $R\_{I6}$ starts near 0 and rises steadily to 300 by step 400, surpassing the others with a robust upward trend. This suggests that $R\_{I6}$ achieves the highest long-term reward, reflecting better adaptability and learning efficiency.
@@ IV-E3 Performance of R_I6 Predictive Model
Our primary focus is on the custom rule $R\_{I6}$, for which we compare the trained value network with the baseline model within the same scenario, featuring a no overtaking sign positioned at 300m along the longitudinal axis. We generate heatmaps using a grid-segmented representation of the road, evaluating the approximated state-value and $R\_{I6}$ rule robustness for both models. As depicted in Figure 5, a yellow grid indicates higher values, while a purple grid signifies lower values. The figure presents a comparative analysis of the trained value network and the baseline model, utilizing heatmaps overlaid on the road grid to assess the approximated state-value and $R\_{I6}$ rule robustness.
The bottom baseline model heatmap displays a sharp yellow-to-purple transition between 250m and 300m, attributable to the traffic sign detection range being set at 50 m. This transition forms a rectangular shape, as the model increasingly avoids leftward positions as it approaches the sign. This suggests that the baseline model relies heavily on immediate rule evaluation, with limited foresight into the no overtaking constraint.
The trained value heatmap exhibits a smooth transition from yellow to purple, approximately aligning with the baseline model. However, the initial value change begins earlier, around 230m, compared to the 250m detection range; the range increased from 50m to around 70m, indicating an awareness of long-term rule violation risks. Notably, the heatmap adopts a trapezoidal shape, with the value in the left lane decreasing earlier than in the right lane. Leveraging the predictive model, the lateral position of the ego vehicle influences the timing of lane-changing maneuvers: a vehicle positioned farther to the left can initiate a rightward lane change earlier, whereas a vehicle already near the right lane may delay the maneuver without violating the rule. This behavior aligns with human intuition and enhances safety and efficiency, as it enables proactive adjustments tailored to the vehicle’s spatial context.
@@ IV-F Discussion
This experiment demonstrates the model’s capability to predict long-term violations of traffic rules within the highD dataset scenarios. The proposed RL approach facilitates safer and more efficient trajectory planning by enabling the system to anticipate potential rule violations and adjust maneuvers proactively, reducing the likelihood of entering states that lead to unavoidable violations. For rules $R\_{I6}$ and $R\_{G1}$, which involve constraints independent of direct vehicle-to-vehicle interactions or focus on maintaining safe distances, the method exhibits strong performance, as evidenced by the smoother transitions in the value network outputs and the expanded safety regions in the heatmaps.
However, the model shows reduced robustness for rules involving complex vehicle interactions, such as $R\_{I2}$, which requires maintaining appropriate speeds relative to surrounding vehicles. This limitation is reflected in the volatility observed in the explained variance metric and the inconsistent value predictions in the heatmap analysis. The challenge arises from the intricate dynamics of vehicle interactions, which demand a more detailed understanding of relative positions, speeds, and potential maneuvers. Addressing this issue may require enhancements such as increasing the volume of training data, extending the training duration, and incorporating additional features, such as relative velocity or lane-changing probabilities, to better capture the interaction dynamics. Furthermore, adopting a more advanced scenario extraction model, such as a Heterogeneous Graph Neural Network, could enhance the model’s ability to represent and manage complex traffic interactions by accounting for diverse vehicles and their interconnected behaviors.
@@ V Conclusion
In this paper, we have introduced an approach to autonomous vehicle path planning that integrates a motion planner with a DRL framework to enhance traffic rule compliance and safety. By replacing the traditional actor network in an AC setup with the cost-based motion planner and leveraging a GNN for state representation, our method effectively combines the interpretability and reliability of motion planning with the predictive power of R. This hybrid framework utilizes the critic’s state-value predictions to guide trajectory selection, enabling the anticipation and prevention of traffic rule violations beyond the immediate planning horizon.
Our approach was evaluated using the highD dataset, focusing on three carefully selected German interstate traffic rules: $R\_G1$, $R\_I2$, and $R\_I6$. Experimental results demonstrate that the proposed method outperforms a non-learning baseline planner, particularly for $R\_I6$ and $R\_G1$, where it achieves smoother value transitions and broader safety margins, as evidenced by heatmap comparisons and metrics such as explained variance and episode reward mean. For $R\_I6$, the model proactively adjusts trajectories to comply with no-overtaking zones, while for $R\_G1$, it anticipates braking events to maintain safe distances. However, performance for $R\_I2$ reveals limitations in handling complex vehicle interactions, exhibiting volatility in predictions that suggests a need for further refinement.
In conclusion, our proposed framework represents a significant step forward in uniting learning-based prediction with explainable planning, offering a scalable and practical solution for autonomous driving. Future work will focus on improving the model’s handling of multi-agent dynamics and exploring its applicability to diverse driving environments beyond German highways.